Grade 12th passPhysical Chemistry

A first order of reaction is complete upto 90%in 10min than calculate the half life period of the reaction

Sourabh
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
Dear student
 
K = 1/t ln (t/ta)
 
K = 1/10 ln 100 /100-90
 
K = (1/10) ln(10) = 0.2303
 
Now
 
As k = 0.693 /t( ½)
 
Hence
 
t( ½) = 0.693 /0.2303 = 3 min
 
Hope it helps